Find the symmetric matrix that corresponds to the quadratic equation 3x12+x1x3 -2x2x3

1
Expert's answer
2021-01-11T17:37:44-0500

We have ternary quadratic form


"f(x_1, x_2, x_3)=3\cdot x_1^2+0\cdot x_1x_2+1\cdot x_1x_3"


"+0\cdot x_2^2-2\cdot x_2x_3+0\cdot x_3^2"

The quadratic form "f(x_1, x_2, x_3)" is determined by the matrix



"A=\begin{bmatrix}
 3 & 0 & 1/2 \\
 0 & 0 & -1 \\
 1/2 & -1 & 0
\end{bmatrix}"